# Keeping Score: A Quantitative Analysis of How the CHI Community Appreciates Its Milestones

### v1

@@ Abstract.
quantitative analysis, milestones, CHI, citations, forgetting curves, bibliometric analysis, Matthew effect, meta-science, meta-HCI.
journalyear: 2024
@@ 1. Introduction
Human-Computer Interaction (HCI) is highly interdisciplinary and characterized by constant evolution and changes (Fuchsberger et al., 2021; Bødker, 2015). Recent changes include, to name but a few, an overall growth in the number of scientific papers published each year (Fortunato et al., 2018; Bornmann et al., 2021), and an acceleration of pace in science with mounting career and publication pressures (Edwards and Roy, 2017; Oppenlaender, 2024). In addition, sudden technological advances can shift focus and topics in HCI research. One such technological advance is the recent progress in large auto-regressive language models, such as ChatGPT (Bubeck et al., 2023; Vaswani et al., 2017). The increasing integration of language models into HCI research reflects broader trends in methodological acceleration and automation (Thomas et al., 2024). While novel AI tools can streamline certain tasks, over-reliance on such tools and their misuse has led to concerns about prompt-hacking (Morris, 2024) and shallow inquiries that prioritize speed over depth (Thomas et al., 2024; Kobak et al., 2024; Gray, 2024; Liang et al., 2024; Academ-AI, 2024; Arawjo, 2024). As HCI evolves, it is critical to reflect on the types of contributions that drive long-lasting impact, particularly amid shifting research norms and topics. In light of this dynamic context, it is important to explore what we value in HCI and understand how the CHI community interacts with its intellectual heritage.
@@
Our research contributes to meta research on HCI literature, particularly focusing on the CHI community, as a not fully representative but important community in the field of HCI.
@@ 3.4. Milestone Coefficient
$\text{SIGN}=\begin{cases}1,&\text{if }m\geq 0\\ -1,&\text{if }m<0\end{cases}\text{with }m=\begin{cases}\text{slope of the linear regression of }c\text{ on }x,&\text{if }n>1\\ 0,&\text{if }n=1\end{cases}$
@@ 4.1.2. Milestone contributions (RQ2)
Comparison of milestone contribution types with existing contribution types in HCI. Contribution Description with examples Wobbrock and Kientz ((2016))* ACM SIGCHI ((2024)$\dagger$ Method A systematic approach or technique used for conducting research or analysis in HCI.
Examples: qualitative methods (e.g., thematic analysis), quantitative methods (e.g., aligned rank transform)
Methodological Methodology
Tool A practical instrument or software designed to perform specific tasks or measurements in HCI studies.
Examples: survey tools (e.g., NASA-TLX, SUS), software libraries (e.g., D3)
Artifact Artifacts
Framework A comprehensive structure outlining key concepts in HCI which organizes knowledge and provides a structured approach for HCI research.
Examples: Feminist HCI, Reflective Design
Theoretical Theory
Agenda A call for future research or practice directions in HCI.
Example: Feminist HCI agenda
Opinion Argument
Guidelines Prescriptive recommendations or best practices for designing or evaluating human-computer interaction.
Example: Guidelines for inter-rater reliability
Methodological Methodology
Model An abstract representation detailing the components and processes of a system or phenomenon in HCI.
Example: Model of personal informatics
Theoretical Theory
Concept A foundational idea or theory that introduces new perspectives or approaches to HCI design and research.
Example: Ambiguity
Theoretical Theory
@@ 4.1.3. Case study: The rise of CHI’s super milestone
In the year 2024, Braun and Clarke’s most-cited milestone (Braun and Clarke, 2006) accounts for about 1 in 500 citations (0.2%) in the CHI proceedings. Given that CHI papers include an average of about 88 references in 2024 (Oppenlaender, 2024), this means we can estimate that, on average, about one in six (5.7) CHI papers cite Braun and Clarke’s work. At present, no other milestone demonstrates such an exceptional performance. In the following section, we examine how the CHI community’s attention to this and other milestones has evolved over time.
@@ 4.2. How has the CHI Community’s Attention to its Milestones evolved over time? (RQ3)
There is, however, a notable shift between the absolute and relative number of citations to milestones in recent years. In absolute terms, citations to the top-300 of CHI’s milestones have increased, including citations to older milestones. The growth in the absolute number of citations in Figure 5 can be attributed to the extreme growth of references in CHI papers since 2016, as observed in prior work (Oppenlaender, 2024). In the 2024 proceedings alone, the CHI community made 3,951 citations to the prior top-300 milestones, compared to only 359 citations ten years earlier in the 2014 proceedings. In relative terms, the 2024 number of citations to milestones pales in comparison to the total number of citations made in that proceedings year (93,639 citations). The top-300 prior milestones only account for a small fraction of citations in 2024 (4.2%). This demonstrates that while the absolute number of citations to milestones has increased, the relative number of citations to milestones has declined. For comparison, the milestones published in the prior five years each received about 3% of all citations in 2002 (top left in Figure 5). In 2024, the relative number of citations to prior milestones is almost ten times lower (bottom right in Figure 5). This highlights that in relative terms, older milestones are being cited less in the CHI community.
@@
We apply the milestone coefficient (MC) to the top-30 milestones (as listed in Table 1). The MC in Table 1 is negative only for one milestone (D3 (Bostock et al., 2011)) which has seen a decline in citations from the CHI community in recent years. All other milestones in the top-30 list are still trending toward growth in citations. To extend this analysis and derive insights on the CHI as a whole, we apply the milestone coefficient to the 3,000 papers most cited by CHI authors. For these papers, the milestone coefficient ranges from -0.0102 to 0.0768. Figure 6 plots the MC for the top 3,000 most cited papers. Each bubble represents one paper in its year of publication. The size of the bubble represents the total number of citations to this milestone. The exceptional super milestone is clearly visible in the year 2006, achieving an MC of 0.0768 (the highest MC in the dataset). The second most cited milestone, the NASA TLX (Hart and Staveland, 1988), is visible in 1988. From Figure 6, it can be derived that the number of milestones has grown over the years. This is unsurprising, given the growth of the CHI proceedings and the growth in citations (Oppenlaender, 2024). However, the figure also demonstrates that many recent milestones struggle to receive the same attention from the CHI community as the past super milestones. Some of the recent milestones even trend towards a decline in citations, and the number of milestones that are being ‘forgotten’ has increased in the last two decades.
@@ 5.1. The CHI Community’s Milestones
Milestones also do not include some other HCI research contribution types, including literature reviews, datasets, and opinions. In particular, we note that literature reviews are a very young contribution type at CHI. Oppenlaender found that literature reviews only became a topic at the CHI Conference after page restrictions were lifted in 2016, but have since grown in popularity (Oppenlaender, 2024). While it has also become more common to cite datasets and code repositories in CHI papers (Oppenlaender, 2024), datasets are not among the top milestones cited at CHI. Opinions are, perhaps, the least known contribution type and, for this reason, not common at CHI.
@@ 5.5. Limitations and Future Work
Fourth, the growth in the number of papers and the increase in the number of references included in CHI papers may contribute to gradual smoothing of the forgetting curves over time. This would not be an issue if changes were proportional (such as a proportional increase in the number of references and papers). However, as described by Oppenlaender, there was a shift in citation behavior in 2016 that constitutes a non-linearity in the CHI community’s trajectory (Oppenlaender, 2024). The forgetting curves should be interpreted with this non-linearity in mind.

### v2

@@ Abstract.
milestones, CHI, citations, forgetting curves, bibliometric analysis, Matthew effect, quantitative analysis, meta-science, meta-HCI.
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713464isbn: 979-8-4007-1394-1/25/04
@@ 1. Introduction
Human-Computer Interaction (HCI) is highly interdisciplinary and characterized by constant evolution and changes (Fuchsberger et al., 2021; Bødker, 2015). Recent changes include, to name but a few, an overall growth in the number of scientific papers published each year (Fortunato et al., 2018; Bornmann et al., 2021) and an acceleration of pace in science with mounting career and publication pressures (Edwards and Roy, 2017; Oppenlaender, 2025). In addition, sudden technological advances can shift focus and topics in HCI research. One such technological advance is the recent progress in large auto-regressive language models, such as OpenAI’s ChatGPT (Bubeck et al., 2023; Vaswani et al., 2017). The increasing integration of language models into HCI research reflects broader trends in methodological acceleration and automation (Thomas et al., 2024). While novel AI tools can streamline certain tasks, over-reliance on such tools and their misuse has led to concerns about prompt-hacking (akin to p-hacking) (Morris, 2024) and shallow inquiries that prioritize speed over depth (Thomas et al., 2024; Kobak et al., 2024; Gray, 2024; Liang et al., 2024; Academ-AI, 2024; Arawjo, 2024). As HCI evolves, it is critical to reflect on the types of contributions that drive long-lasting impact, particularly amid shifting research norms and topics. In light of this dynamic context, it is important to explore what we value in HCI and to understand how the CHI community interacts with its intellectual heritage.
@@
Our research contributes to meta research on HCI literature (“meta-HCI”) (Oppenlaender et al., 2025), particularly focusing on the CHI community, as a not fully representative but important community in the field of HCI.
@@ 3.4. Milestone Coefficient
$\text{SIGN}=\begin{cases}1,&\text{if }m\geq 0\\ -1,&\text{if }m<0\end{cases}$
$m=\begin{cases}\text{slope of the linear regression of }c\text{ on }x,&\text{if }n>1\\ 0,&\text{if }n=1\end{cases}$
@@ 4.1.2. Milestone contributions (RQ2)
Comparison of milestone contribution types with existing contribution types in HCI. Contribution Description with examples Wobbrock and Kientz ((2016))* ACM SIGCHI ((2024)$\dagger$ Method A systematic approach or technique used for conducting research or analysis in HCI.Examples: qualitative methods (e.g., thematic analysis), quantitative methods (e.g., aligned rank transform) Methodological Methodology Tool A practical instrument or software designed to perform specific tasks or measurements in HCI studies. Examples: survey tools (e.g., NASA-TLX, SUS), software libraries (e.g., D3) Artifact Artifacts Framework A comprehensive structure outlining key concepts in HCI which organizes knowledge and provides a structured approach for HCI research. Examples: Feminist HCI, Reflective Design Theoretical Theory Agenda A call for future research or practice directions in HCI.Example: Feminist HCI agenda Opinion Argument Guidelines Prescriptive recommendations or best practices for designing or evaluating human-computer interaction. Example: Guidelines for inter-rater reliability Methodological Methodology Model An abstract representation detailing the components and processes of a system or phenomenon in HCI. Example: Model of personal informatics Theoretical Theory Concept A foundational idea or theory that introduces new perspectives or approaches to HCI design and research.Example: Ambiguity Theoretical Theory
@@ 4.1.3. Case study: The rise of CHI’s super milestone
In the year 2024, Braun and Clarke’s most-cited milestone (Braun and Clarke, 2006) accounts for about 1 in 500 citations (0.2%) in the CHI proceedings. Given that CHI papers include an average of about 88 references in 2024 (Oppenlaender, 2025), this means we can estimate that, on average, about one in six (5.7) CHI papers cite Braun and Clarke’s work. At present, no other milestone demonstrates such an exceptional performance. In the following section, we examine how the CHI community’s attention to this and other milestones has evolved over time.
@@ 4.2. How has the CHI Community’s Attention to its Milestones evolved over time? (RQ3)
There is, however, a notable shift between the absolute and relative number of citations to milestones in recent years. In absolute terms, citations to the top-300 of CHI’s milestones have increased, including citations to older milestones. The growth in the absolute number of citations in Figure 5 can be attributed to the extreme growth of references in CHI papers since 2016, as observed in prior work (Oppenlaender, 2025). In the 2024 proceedings alone, the CHI community made 3,951 citations to the prior top-300 milestones, compared to only 359 citations ten years earlier in the 2014 proceedings. In relative terms, the 2024 number of citations to milestones pales in comparison to the total number of citations made in that proceedings year (93,639 citations). The top-300 prior milestones only account for a small fraction of citations in 2024 (4.2%). This demonstrates that while the absolute number of citations to milestones has increased, the relative number of citations to milestones has declined. For comparison, the milestones published in the prior five years each received about 3% of all citations in 2002 (top left in Figure 5). In 2024, the relative number of citations to prior milestones is almost ten times lower (bottom right in Figure 5). This highlights that in relative terms, older milestones are being cited less in the CHI community.
@@
We apply the milestone coefficient (MC) to the top-30 milestones (as listed in Table 1). The MC in Table 1 is negative only for one milestone (D3 (Bostock et al., 2011)) which has seen a decline in citations from the CHI community in recent years. All other milestones in the top-30 list are still trending toward growth in citations. To extend this analysis and derive insights on the CHI as a whole, we apply the milestone coefficient to the 3,000 papers most cited by CHI authors. For these papers, the milestone coefficient ranges from -0.0102 to 0.0768. Figure 6 plots the MC for the top 3,000 most cited papers. Each bubble represents one paper in its year of publication. The size of the bubble represents the total number of citations to this milestone. The exceptional super milestone is clearly visible in the year 2006, achieving an MC of 0.0768 (the highest MC in the dataset). The second most cited milestone, the NASA TLX (Hart and Staveland, 1988), is visible in 1988. From Figure 6, it can be derived that the number of milestones has grown over the years. This is unsurprising, given the growth of the CHI proceedings and the growth in citations (Oppenlaender, 2025). However, the figure also demonstrates that many recent milestones struggle to receive the same attention from the CHI community as the past super milestones. Some of the recent milestones even trend towards a decline in citations, and the number of milestones that are being ‘forgotten’ has increased in the last two decades.
@@ 5.1. The CHI Community’s Milestones
Milestones also do not include some other HCI research contribution types, including literature reviews, datasets, and opinions. In particular, we note that literature reviews are a very young contribution type at CHI. Oppenlaender found that literature reviews only became a topic at the CHI Conference after page restrictions were lifted in 2016, but have since grown in popularity (Oppenlaender, 2025). While it has also become more common to cite datasets and code repositories in CHI papers (Oppenlaender, 2025), datasets are not among the top milestones cited at CHI. Opinions are, perhaps, the least known contribution type and, for this reason, not common at CHI.
@@ 5.5. Limitations and Future Work
Fourth, the growth in the number of papers and the increase in the number of references included in CHI papers may contribute to gradual smoothing of the forgetting curves over time. This would not be an issue if changes were proportional (such as a proportional increase in the number of references and papers). However, as described by Oppenlaender, there was a shift in citation behavior in 2016 that constitutes a non-linearity in the CHI community’s trajectory (Oppenlaender, 2025). The forgetting curves should be interpreted with this non-linearity in mind.
@@ 6. Conclusion
Acknowledgements.
This research was partially funded by the Strategic Research Council (SRC), established within the Academy of Finland (Grants 356128, 335625, 335729), and Academy Research Fellow funding by Academy of Finland (Grants 356128, 349637 and 353790).